In 1996, NASA performed an experiment called the Tethered Satellite experiment. In this experiment a 1.60 104-m length of wire was let out by the space shuttle Atlantis to generate a motional emf. The shuttle had an orbital speed of 7.60 103 m/s, and the magnitude of the earth's magnetic field at the location of the wire was 5.10 10-5 T. If the wire had moved perpendicular to the earth's magnetic field, what would have been the motional emf generated between the ends of the wire?
Answer #1

ε = B * L * v

ε = 5*10-5 * 1.6*104 *7.6*103 = 6080V

I hope it was helpful.. please rate me positive